In how many ways can the committee of 4 women and 2 men can be selected from a group of 10 women and 5 men If there are no restriction?

1
Expert's answer
2021-12-01T18:12:53-0500
"C(10, 4)C(5, 2)=\dbinom{10}{4}\dbinom{5}{2}"

"=\dfrac{10!}{4!(10-4)!}\cdot\dfrac{5!}{2!(5-2)!}=210\cdot10"


"=2100"

The committee of 4 women and 2 men can be selected from a group of 10 women and 5 men If there are no restriction, in 2100 ways.